OpenStudy (anonymous):

Consider the curve given by x^2+4y^2=7+3xy show that dy/dx=(3y-2x)/(8y-3x) how do you differentiate the 3xy part I'm confused?

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( 3 x y \right)=3\left( x \frac{ dy }{ dx }+y \frac{ d }{ dx } \left( x \right)\right)\]